OpenStudy (anonymous):

Draw the exponential which passes through the points (-1, 8) and (3, 3). From your graph, estimate the halving time from your graph. Find a formula for the exponential, and estimate the halving time from your formula by guessing and checking. Using logarithms, nd the exact halving time and compare this to your estimate.

OpenStudy (anonymous):

So far by doing y=m*x+b and using \[\frac{ 3-8 }{ 3-(-1) } = \frac{ -5 }{4} = m\]

OpenStudy (anonymous):

Then working out out: 8 = -5/4*(-1)+b = 5/4 + 27/4 Thus m = -5/4 b = 27/4 Is this all corret so far?

OpenStudy (anonymous):

Just found also using y= a(b^x), Which I assume is how I proceed?

OpenStudy (anonymous):

Basically, you have two equations and two variables: \[ 8=ab^{-1} \\ 3=ab^3 \]

OpenStudy (anonymous):

Solve for \(a\) in terms of \(b\), then substitute it into the other equation.

OpenStudy (anonymous):

If I multiply 8*b^3 and 3*b^-1 will this help find the b value?

OpenStudy (anonymous):

I am using this site http://www.ehow.com/how_8117999_exponential-equation-two-points.html but it is confusing on how to proceed.

OpenStudy (anonymous):

Okay first of all in the first equation we solve for \(a\) \[ 8=ab^{-1}\implies 8b=a \]Then we substitute in to the second equation \[ 3 = ab^3 = (8b)b^3 = 8b^4 \]

OpenStudy (anonymous):

This means \[ b = \sqrt[4]{\frac{3}{8}} \]

OpenStudy (anonymous):

@Astrobuoy Do you see what I did there?

OpenStudy (anonymous):

thus \[8*\sqrt[4]{\frac{ 3 }{ 8 }} = a\] Hence: \[4*\sqrt[4]{6} = a\]

OpenStudy (anonymous):

Hmm, now did you get that?

OpenStudy (anonymous):

\[ 8\sqrt[4]{\frac{3}{8}}=\sqrt[4]{\frac{8^4}{1}}\sqrt[4]{\frac{3}{8}} = \sqrt[4]{3\cdot 8^3} \]

OpenStudy (anonymous):

You wrote eariler 8b = a, So I mutiplied 8 by the b that is listed there.

OpenStudy (anonymous):

Strange, I did the problem on Wolfram and it came up as what I wrote. :\

OpenStudy (anonymous):

It is strange.

OpenStudy (anonymous):

Hmm, maybe they did \(8=4\cdot 2=4\sqrt[4]{16}\)

OpenStudy (anonymous):

Perhaps a syntax error on mine or the systems behalf. So my final a = \[\sqrt[4]{3*8^{3}} ?\]

OpenStudy (anonymous):

no, I think actually it is correct, I just didn't simply mine enough.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=8*%283%2F8%29%5E%281%2F4%29%29 Is the link from my page of computation.

OpenStudy (anonymous):

anyway we we done?

OpenStudy (anonymous):

cheers, I can take it from here :) thanks for the help!